# Acoustic-to-articulatory inversion for dysarthric speech: Are pre-trained self-supervised representations favorable?

### 22 Dec 2023 v3

@@ 1 Introduction
Cerebral Palsy (CP) is a neurological disorder that affects movement, balance, and posture, often caused by damage to the developing brain before or during birth. It is a non-progressive condition. Amyotrophic Lateral Sclerosis (ALS), however, is a progressive neurological disease that leads to muscle weakness and eventual paralysis. Due to a slow muscle response time of a subject, the speech-motor functions are severely affected [1], which eventually leads to weak pronunciation, an affected accent, and unintelligible speech, causing dysarthria [2]. As the severity of dysarthria increases, it has a negative impact on the movement of articulators, such as the lips, jaw, tongue, and velum, which are responsible for speech production [3, 4].
@@ 2 Dataset
In this work, we use the TORGO dataset [22] that comprises aligned speech acoustics and measured 3D EMA articulatory trajectories from speakers with either CP or ALS and matched healthy controls. We choose only 4 speakers with complete parallel acoustic-articulatory data i.e., 2 male healthy controls (MC01 and MC04) and 2 female patients (F03 and F04). Articulatory movements, in the X and Y directions, from sensor coils attached to the tongue tip (TT), tongue middle (TM), tongue back (TB), jaw (JAW), lower lip (LL), and upper lip (UL) are considered. The tongue’s middle and back portions are also called the body and dorsum, respectively. We denote TM and TB as TB and TD, henceforth. This results in 12-dimensional articulatory feature vectors denoted as UL, UL, LL, LL, JAW, JAW, TT, TT, TB, TB, TD, and TD. Illa et al. [3] proposed the addition of kinematic features like velocity and acceleration to the articulatory features in order to improve speech-based classification between ALS patients and healthy controls. This results in a 6-dimensional velocity and acceleration articulatory feature vector along the X and Y directions, each. Finally, we obtain 24-dimensional articulatory feature vectors upon concatentation.
@@ 4 Experimental Setup
Feature extraction: As an initial step, we pre-process the recorded speech waveforms and their corresponding articulatory features. The speech waveforms are downsampled to 16 kHz. The articulatory features are downsampled to 100 Hz, from the initial 200 Hz, and low-pass filtered at a cutoff frequency of 25 Hz. This was done due to the presence of high-frequency noise that could be incurred during the recording session. We use the Kaldi toolkit [27] to compute MFCCs, using a window length of 25ms and a shift of 10ms. This way, the MFCCs have a one-to-one correspondence with the articulatory trajectories. For SSL feature extraction, we use the upstream model weights from s3prl 111https://github.com/s3prl/s3prl, for the initialization of each of the SSL models. Mean and variance normalization is performed at an utterance level, across each dimension of the acoustic and articulatory features. The SSL feature representations are derived from the final layer of the SSL model [10]. For the computation of x-vectors, we use the Kaldi toolkit [27], using a pre-trained model trained on the VoxCeleb database [28]. Fig. 2 illustrates the x-vector speaker embeddings, for each subject, after t-SNE [29]. We observe that the embeddings obtained from the pre-trained model are able to discriminate between the speakers. However, an interesting observation is that the x-vectors for certain sentences from F03 overlap with those of healthy controls. This could be due to the low dysarthric severity of F03 which translates to similar speaking rates or styles. In addition, we also observe that the x-vectors of MC01 form two separate clusters. This could be due to different speaking rates and lexical patterns across sentences for MC01.
@@
Training schemes and network parameters: We perform experiments in two subject conditions - seen and unseen. In the seen conditions, for each subject, we uniformly split the sentences as 90% for training and the remaining 10% for testing. We make sure that the sentences do not overlap across the two sets. Here, the AAI models are trained in three training schemes: • Subject-specific: AAI model is trained on the training data of each subject and tested on its corresponding test set. • Pooled: A single AAI model is trained on a combined set of training data of all subjects. During inference, the test sets of the subjects are aggregated. • Fine-tuned: Trained parameters from the pooled scheme are initialized and fine-tuned on the training data of each subject and re-trained. In the unseen subject conditions, we employ the leave-one-person-out approach i.e., we train an AAI model on the data of all subjects except the one chosen as the test subject. We test the AAI model on the subject that is left out. In both the training conditions (seen and unseen), we consider a standard 5-fold cross-validation setup out of which 4 folds are considered for training and the other fold is considered for validation, in a round-robin fashion.
@@ 5.2 Impact of different pre-trained SSL features
Seen subject evaluation: Table 3 reports the average CC (standard deviation), in the seen subject conditions, across the respective test sets, averaged across all the articulators, sentences, and folds, for all the training schemes i.e., subject-specific, pooled, and fine-tuned, for the different SSL features, in this work, as the input. In the subject-specific scheme, we notice that features from wav2vec and DeCoAR perform better over MFCCs, for both healthy controls and patients. In particular, DeCoAR performs the best with relative CC improvements of 0.94% and 5.34% over MFCCs, for healthy controls and patients respectively. It is interesting to note that vq_wav2vec, Mockingjay, and TERA, perform poorly, but APC and NPC achieve comparable performances to MFCCs. The drop in performance by vq_wav2vec could be due to the quantization step in its training objective [12]. TERA and Mockingjay which are trained with a masked loss function to predict the sequence of words in a sentence, seem to give a lower performance [13, 18] over here. In addition, a stronger reason could be that in the subject-specific scheme, the amount of training data from each subject is less, contributing to overfitting and poor generalization on the corresponding test set. For the pooled and fine-tuned training schemes, we observe that the performances of all the SSL features are better than MFCCs, for both healthy control and patients, except for Mockingjay and vq_wav2vec. However, vq_wav2vec performs better than MFCCs for patients. To be specific, DeCoAR performs the best across all the setups with CC improvements of 1.81% and 4.56% for healthy controls and patients respectively, in the fine-tuned case. The combination of unsupervised pre-training and supervised fine-tuning, in DeCoAR, allows the network to learn general features of speech and task-specific features [16] i.e., articulatory trajectories of healthy controls and patients. This validates that SSL pre-trained features are more suited and robust acoustic features for dysarthric AAI.
Unseen subject evaluation: Table 4 reports the average CC (standard deviation), in the unseen subject conditions, across all the sentences and articulators, for each subject. We observe similar average performances, as in the seen conditions, across the different subjects. In particular, wav2vec and DeCoAR perform better over MFCCs, for all the subjects. For F03 (with moderate dysarthric severity), the CC can be improved from 0.4201 to 0.4659, using DeCoAR. It is to be noted that AAI on unseen dysarthric subjects is challenging. Pre-trained SSL features combined with the conditioning of x-vectors, led to a better utterance level generalisability on subjects unseen during training [31]. This confirms the effectiveness of pre-trained SSL features in performing unseen dysarthric AAI.
@@
Effect of data adaptation: Here, we study the amount of training data required from an unseen target subject to achieve a comparable performance on using the same subject’s complete training data. We begin by selecting t% of training data from a subject and fine-tune the AAI model weights that have been trained on the remaining reference subjects. In all the plots of Fig. 3, the x-axis reports the various t% of training data that we choose and the y-axis reports the average CC across all the articulators, sentences, and folds. Clearly, t=0 refers to an unseen subject evaluation. From Fig. 3, we observe that for all the subjects with t0, DeCoAR performs better than MFCCs. The better performances of DeCoAR by using small amounts of training data from a target subject could be attributed to the rich representation space and multiple acoustic-articulatory mappings it learns, for the dysarthric patients and healthy controls. For F03 and F04 at the t=50 mark, DeCoAR archives relative improvements of 11.23% and 6.59% respectively, over MFCCs. Overall, experimental results suggest that using small amounts of training data (instead of full training, t=100) with pre-trained SSL features, from an unseen target subject, is beneficial for dysarthric AAI since collecting full acoustic-articulatory data from a subject is tedious and time-consuming.
@@ 6 Conclusions
We introduce an approach that uses pre-trained SSL features for AAI on dysarthric speech and demonstrate their effects. We perform experiments in seen and unseen subject conditions. The results revealed that pre-trained SSL features (with x-vectors) are beneficial in learning the AAI mapping for dysarthric patients. In particular, features from wav2vec, DeCoAR, and APC perform well or are on par with baseline MFCCs for dysarthric AAI in both seen and unseen conditions. We also showed that using smaller amounts of training data from an unseen target subject and trained using DeCoAR, as the input acoustic feature, is crucial to achieve a superior performance. Our future work includes understanding the effects of SSL networks on a language-mismatched dysarthric corpus with high severity.
7 Acknowledgements
We would like to acknowledge Dr. Cauligi S. Raghavendra for his valuable comments and support throughout.

### 09 Feb 2024 v4

@@ 1 Introduction
Cerebral Palsy (CP) is a neurological disorder that affects movement, balance, and posture, often caused by damage to the developing brain before or during birth. It is a non-progressive condition. Amyotrophic Lateral Sclerosis (ALS), however, is a progressive neurological disease that leads to muscle weakness and eventual paralysis. Due to a slow muscle response time of a subject, the speech-motor functions are severely affected [1], which eventually leads to weak pronunciation, an affected accent, and unintelligible speech, causing dysarthria [2]. As the severity of dysarthria increases, it hurts the movement of articulators, such as the lips, jaw, tongue, and velum, which are responsible for speech production [3, 4].
@@ 2 Dataset
In this work, we use the TORGO dataset [22] that comprises aligned speech acoustics and measured 3D EMA articulatory trajectories from 15 speakers with either CP or ALS and matched healthy controls. Out of the acoustic-articulatory data available from 7 speakers only, we choose 4 speakers with complete parallel acoustic-articulatory data i.e., 2 male healthy controls (MC01 and MC04) and 2 female patients (F03 and F04). Articulatory movements, in the X and Y directions, from sensor coils attached to the tongue tip (TT), tongue middle (TM), tongue back (TB), jaw (JAW), lower lip (LL), and upper lip (UL) are considered. The tongue’s middle and back portions are also called the body and dorsum, respectively. We denote TM and TB as TB and TD, henceforth. This results in 12-dimensional articulatory feature vectors denoted as UL, UL, LL, LL, JAW, JAW, TT, TT, TB, TB, TD, and TD. Illa et al. [3] proposed the addition of kinematic features like velocity and acceleration to the articulatory features to improve speech-based classification between ALS patients and healthy controls. This results in a 6-dimensional velocity and acceleration articulatory feature vector along the X and Y directions, each. Finally, we obtain 24-dimensional articulatory feature vectors upon concatenation.
@@ 4 Experimental Setup
Feature extraction: As an initial step, we pre-process the recorded speech waveforms and their corresponding articulatory features. The speech waveforms are downsampled to 16 kHz. The articulatory features are downsampled to 100 Hz, from the initial 200 Hz, and low-pass filtered at a cutoff frequency of 25 Hz. This was done due to the presence of high-frequency noise that could be incurred during the recording session. We use the Kaldi toolkit [27] to compute MFCCs, using a window length of 25ms and a shift of 10ms. This way, the MFCCs have a one-to-one correspondence with the articulatory trajectories. For SSL feature extraction, we use the upstream model weights from s3prl 111https://github.com/s3prl/s3prl, for the initialization of each of the SSL models. We initialize using the default pre-trained weights, as provided by s3prl. Mean and variance normalization is performed at an utterance level, across each dimension of the acoustic and articulatory features. The SSL feature representations are derived from the final layer of the SSL model [10]. For the computation of x-vectors, we use the Kaldi toolkit [27], using a pre-trained model trained on the VoxCeleb database [28]. Fig. 2 illustrates the x-vector speaker embeddings, for each subject, after t-SNE [29]. We observe that the embeddings obtained from the pre-trained model are able to discriminate between the speakers. However, an interesting observation is that the x-vectors for certain sentences from F03 overlap with those of healthy controls. This could be due to the low dysarthric severity of F03 which translates to similar speaking rates or styles. In addition, we also observe that the x-vectors of MC01 form two separate clusters. This could be due to different speaking rates and lexical patterns across sentences for MC01.
@@
Training schemes and network parameters: We perform experiments in two subject conditions - seen and unseen. In the seen conditions, for each subject, we uniformly split the sentences as 90% for training and the remaining 10% for testing. We make sure that the sentences do not overlap across the two sets. Here, the AAI models are trained in three training schemes: • Subject-specific: The AAI model is trained on the training data of each subject and tested on its corresponding test set. • Pooled: A single AAI model is trained on a combined set of training data of all subjects. During inference, the test sets of the subjects are aggregated. • Fine-tuned: Trained parameters from the pooled scheme are initialized and fine-tuned on the training data of each subject and re-trained. In the unseen subject conditions, we employ the leave-one-person-out approach i.e., we train an AAI model on the data of all subjects except the one chosen as the test subject. We test the AAI model on the subject that is left out. In both the training conditions (seen and unseen), we consider a standard 5-fold cross-validation setup out of which 4 folds are considered for training and the other fold is considered for validation, in a round-robin fashion.
@@ 5.2 Impact of different pre-trained SSL features
Seen subject evaluation: Table 3 reports the average CC (standard deviation), in the seen subject conditions, across the respective test sets, averaged across all the articulators, sentences, and folds, for all the training schemes i.e., subject-specific, pooled, and fine-tuned, for the different SSL features, in this work, as the input. In the subject-specific scheme, we notice that features from DeCoAR perform better over MFCCs with relative CC improvements of 0.94% and 5.34% over MFCCs, for healthy controls and patients respectively. It is interesting to note that vq_wav2vec, Mockingjay, and TERA, perform marginally poorly, but APC and NPC achieve comparable performances to MFCCs. The drop in performance by vq_wav2vec could be due to the quantization step in its training objective [12]. TERA and Mockingjay which are trained with a masked loss function to predict the sequence of words in a sentence, seem to give a lower performance [13, 18] over here. In addition, a stronger reason could be that in the subject-specific scheme, the amount of training data from each subject is less, contributing to overfitting and poor generalization on the corresponding test set. For the pooled and fine-tuned training schemes, we observe that the performances of all the SSL features are better than MFCCs, for both healthy control and patients, except for Mockingjay and vq_wav2vec. However, vq_wav2vec performs better than MFCCs for patients. To be specific, DeCoAR performs the best across all the setups with CC improvements of 1.81% and 4.56% for healthy controls and patients respectively, in the fine-tuned case. The combination of unsupervised pre-training and supervised fine-tuning, in DeCoAR, allows the network to learn general features of speech and task-specific features [16] i.e., articulatory trajectories of healthy controls and patients. This validates that SSL pre-trained features are more suited and robust acoustic features for dysarthric AAI. Interestingly, on the whole, the results using MFCCs for patients are competitive due to lower dysarthria severity.
Unseen subject evaluation: Table 4 reports the average CC (standard deviation), in the unseen subject conditions, across all the sentences and articulators, for each subject. We observe similar average performances, as in the seen conditions, across the different subjects. In particular, wav2vec and DeCoAR perform better over MFCCs, for all the subjects. For F03 (with moderate dysarthric severity), the CC can be improved from 0.4201 to 0.4659, using DeCoAR. It is to be noted that AAI on unseen dysarthric subjects is challenging. Pre-trained SSL features combined with the conditioning of x-vectors led to a better utterance level generalisability on subjects unseen during training [31]. This confirms the effectiveness of pre-trained SSL features in performing unseen dysarthric AAI.
@@
Effect of data adaptation: Here, we study the amount of training data required from an unseen target subject to achieve a comparable performance using the same subject’s complete training data. We begin by selecting t% of training data from a subject and fine-tuning the AAI model weights that have been trained on the remaining reference subjects. In all the plots of Fig. 3, the x-axis reports the various t% of training data that we choose, and the y-axis reports the average CC across all the articulators, sentences, and folds. Clearly, t=0 refers to an unseen subject evaluation. From Fig. 3, we observe that for all the subjects with t0, DeCoAR performs better than MFCCs. The better performances of DeCoAR by using small amounts of training data from a target subject could be attributed to the rich representation space and multiple acoustic-articulatory mappings it learns, for the dysarthric patients and healthy controls. For F03 and F04 at the t=50 mark, DeCoAR archives relative improvements of 11.23% and 6.59% respectively, over MFCCs. Overall, experimental results suggest that using small amounts of training data (instead of full training, t=100) with pre-trained SSL features, from an unseen target subject, is beneficial for dysarthric AAI since collecting full acoustic-articulatory data from a subject is tedious.
@@ 6 Conclusions
We introduce an approach that uses pre-trained SSL features, with x-vectors, for AAI on dysarthric speech and demonstrate their effects. We perform experiments in seen and unseen subject conditions. The results revealed that, in particular, features from wav2vec, DeCoAR, and APC perform well or are on par with baseline MFCCs for dysarthric AAI in both seen and unseen conditions. We also showed that using smaller amounts of training data from an unseen target subject and trained using DeCoAR, as the input acoustic feature, achieves a superior performance owing to its unsupervised pre-training and supervised fine-tuning training scheme. Our future work includes understanding the effects of SSL networks on a language-mismatched dysarthric corpus with high severity.